# Beyond Memorization: A Rigorous Evaluation Framework for Medical Knowledge Editing

### v2

@@ Abstract
Recently, knowledge editing (KE) has emerged as a promising approach to update specific facts in Large Language Models (LLMs) without the need for full retraining. Despite the effectiveness in general-domain benchmarks, their applicability to complex medical domain remains largely unexplored. Medical knowledge editing is particularly challenging, as it requires LLMs to internalize the knowledge and generalize to unseen scenarios for effective and interpretable decision-making. In this work, we propose a novel framework called MedEditBench to rigorously evaluate the effectiveness of existing KE methods in the medical domain. In MedEditBench, we introduce a new medical knowledge editing benchmark as well as three different knowledge editing paradigms, which are designed to assess the impact of different knowledge sources for editing. Our findings indicate that current KE methods result in only superficial memorization of the injected information, failing to generalize to new scenarios. To overcome this limitation, we present Self-Generated Rationale Editing (SGR-Edit), which utilizes model-derived rationales as the target knowledge for editing, thereby uncovering the underlying reasoning process and demonstrating significant improvements over existing KE approaches. Additionally, we offer deeper insights into medical knowledge editing, including the localization of medical knowledge in LLMs and the impact of sequential editing on evolving knowledge. This could provide practical guidance for implementing KE methods in real-world medical applications.  [^1] Codes and the MedEditBench datasets are released at: https://github.com/Aries-chen/MedEditBench
@@
Shigeng Chen [^1] footnotemark: 1,  Linhao Luo1thanks:   Equal Contribution. <arieschen0224@gmail.com>.,  Zhangchi Qiu1, Yanan Cao2,  Carl Yang3,  Shirui Pan1thanks:   Corresponding Author. <s.pan@griffith.edu.au > 1Griffith University,  2Chinese Academy of Sciences, Beijing, China, 3Emory University
@@ 1 Introduction
Large language models (LLMs) encapsulate extensive knowledge during training on large-scale corpora (Petroni et al., 2019; Allen-Zhu and Li, 2023; Dai et al., 2021). Nonetheless, their knowledge remains static after training, resulting in factual inconsistencies and hallucinations in tasks that require up-to-date information or domain-specific expertise beyond their pre-trained knowledge (Hu et al., 2023; Huang et al., 2025). Knowledge editing (KE) has emerged as a promising approach to update specific knowledge in LLMs without full retraining (Zhang et al., 2024; Wang et al., 2024b), which provides an efficient and effective solution to constantly adapt LLMs to the ever-evolving world of knowledge (Meng et al., 2022, ; Fang et al., 2024).
@@
Medical knowledge editing is particularly challenging due to its intricacy and specificity, which requires LLMs to not only memorize updated medical knowledge (Singhal et al., 2023) but also comprehend underlying medical concepts and generalize to new, unseen scenarios for effective reasoning (Li et al., 2023a), as shown in Figure 1. Moreover, the high stakes of medical applications necessitate that LLMs explain their reasoning to improve the trustfulness and interoperability (Holzinger et al., 2017; Tonekaboni et al., 2019). Therefore, it is crucial to establish a more rigorous evaluation framework that accurately analyzes the performance of existing KE methods on medical knowledge to ensure their reliability and applicability in real-world applications (Zhou et al., 2024; Zhang et al., 2025).
In this work, we establish MedEditBench for rigorous medical knowledge editing evaluation. First, we create a high-quality medical knowledge editing benchmark from two real-world medical question-answering datasets (Kim et al., 2024; Pal et al., 2022). To evaluate the generalizability of KE methods, we extend the benchmark with scenario-based questions that require LLMs to apply the injected knowledge to address new clinical scenarios, as well as questions that require LLMs to preserve their previously factual knowledge. We also introduce three novel metrics: efficacy, generalization, and retention to measure the performance of KE methods in updating, generalizing, and preserving medical knowledge, respectively.
Our initial findings reveal a substantial gap between current KE methods and the demands of real-world medical applications. We observe that all existing KE methods consistently underperform in complex medical settings. Even the strongest method, AlphaEdit (Fang et al., 2024), achieves only a post-edit score of 53.9% (the mean of 43.9% efficacy, 31.2% generalization, and 86.7% retention). We attribute this pitfall to the existing editing paradigm that relies on a short ground-truth answer (GTA) as the knowledge target, which causes LLMs to memorize surface-level facts rather than understand the underlying medical rationale, thereby impairing generalization. Although using the human-curated reference (RE) as the target knowledge can enhance the knowledge understanding by providing more context, the improvements are still limited.
To address this limitation, we introduce a novel editing paradigm called self-generated rationale editing (SGR-Edit), which simulates the chain-of-thought reasoning (Wei et al., 2022), by prompting LLMs to generate explanatory rationales to support the new target answer based on provided reference texts. These self-generated rationales then serve as the target knowledge and can be seamlessly integrated with arbitrary KE methods. SGR-Edit reveals the internal reasoning process of LLMs, allowing them to better internalize new medical knowledge and generalize to unfamiliar scenarios. Experimental results demonstrate that SGR-Edit significantly improves the performance of existing KE methods compared to the current paradigm of editing with a short GTA. Overall, SGR-Edit yields an average absolute improvement of 8.6 percentage points over GTA-Edit on LLaMA-8B and 12.6 percentage points on LLaMA-3B, demonstrating its effectiveness in enhancing medical knowledge editing. The contribution of this work is as follows.
  • We propose a framework (MedEditBench) to rigorously assess existing knowledge editing methods in real-world medical scenarios and evaluate the applicability of various editing paradigms with different target knowledge.
  • We conduct a comprehensive evaluation and find that existing editing methods are insufficient for medical knowledge editing because the existing paradigm relies on short answers, promoting surface-level memorization.
  • We propose a novel editing paradigm, SGR-Edit, which uses self-generated rationales as the target knowledge for editing, noticeably improving the performance of all existing editing methods.
  • We introduce deeper insights into medical knowledge editing, including the localization of medical knowledge in LLMs and the impact of sequential editing on evolving knowledge.
@@ 2.1 Knowledge Editing Methods
Knowledge editing methods can be roughly grouped into three categories:
Fine-Tuning–Based Editing
These approaches update a large number of model parameters to inject new knowledge, typically via constrained or parameter-efficient training. Such approaches include FT+L (Zhu et al., 2020), FT-M (Zhang et al., 2024), and LoRA (Hu et al., 2022).
Parameter-Modifying Editing
Compared to full model fine-tuning, these methods focus on specific parameters to minimize interference with unrelated knowledge. The Meta-Learning strategy trains a hypernetwork to predict gradient updates for knowledge insertion, as seen in methods (De Cao et al., 2021; Mitchell et al., ; Tan et al., ). The Locate-then-Edit approach identifies and rewrites factual weights in specific layers. Representative models of this approach include ROME (Meng et al., 2022) and MEMIT (Meng et al., ), along with sequential editing variants like PRUNE (Ma et al., 2024a), AlphaEdit (Fang et al., 2024), and AnyEdit (Jiang et al., 2025), which support continuous knowledge updates.
Parameter-Preserving Editing
These techniques maintain the base model’s parameters by either augmenting it with external modules or retrieving relevant information during inference. Extension-Based methods, such as GRACE (Hartvigsen et al., 2023), add adapters or incorporate side modules (Mitchell et al., 2022; Wang et al., 2024a) to store new facts. Retrieval-Based methods (Zheng et al., 2023; Song et al., 2024; Shi et al., 2024; Chen et al., 2024b) retrieve relevant facts to include in the prompt as context and generate updated outputs.
@@ 2.2 Knowledge Editing Benchmarks
Early evaluations of knowledge editing methods primarily focused on general-domain benchmarks, such as WikiData (Meng et al., 2022) and counterfactual datasets (Levy et al., 2017). They try to update LLMs with new knowledge that contradicts the common knowledge. Recently, studies (Ju et al., 2023; Li et al., 2023b; Pinter and Elhadad, 2023; Chen et al., 2024a) question the real-world applicability of existing knowledge editing methods and propose new benchmarks for fair evaluations.
Recently, Cohen et al. (2024) measures “ripple effects” on related facts and reveals that most KE methods often fail to propagate consistent changes beyond the target triple. Huang et al. (2024) reports that prior benchmarks do not strictly confirm LLMs having hallucinated answers to the questions before conducting editing, which masks their true editing performance. Ma et al. (2024b) investigates editing consistency under prompt rephrasing and realistic communicative contexts, finding that current KEs exhibit lower generalization and that popular facts are hardest to edit. Lin et al. (2024) examines sequential editing and shows that KE methods experience editing performance drops after continual edits. Yang et al. (2025) critiques common evaluation practices and reveals that KE methods catastrophically fail on realistic QA tasks. Differ from prior evaluations focused on artificial and simplified general-domain settings (Ma et al., 2024b; Lin et al., 2024), we introduce the first framework for medical knowledge editing evaluation, to rigorously measure whether updated medical knowledge could be effectively adopted in new clinical scenarios without degrading previous factual knowledge.
@@ 3 Task Formulation
The goal of knowledge editing is to modify specific knowledge $k$ in LLMs without retraining the entire model (Zhang et al., 2024), thereby improving performance on tasks related to that knowledge, represented by a set of queries and answers: $\mathcal{Q}^{k}=\{(q_{i},a_{i})\}$. Let $\theta$ denote the original model. Given a query $q$ and the target knowledge $k$, the knowledge editing method $F$ can be expressed as:
@@ 4 Evaluation Framework
Figure 2: Overview of proposed medical knowledge editing evaluation framework (MedEditBench). Datasets are first constructed for medical knowledge editing (gray rectangle). Then, we propose three editing paradigms (GTA-Edit, RE-Edit, and SGR-Edit) to evaluate the effectiveness of different target knowledge for medical knowledge editing (purple rectangle). The rationale in SGR-Edit is generated by the LLM itself, given a QA and a reference (green rectangle). These evaluations are based on the existing knowledge editing methods (blue rectangle). Finally, we evaluate the edited model’s post-edit accuracy (i.e., efficacy, generalization, and retention) and assess the interpretability of the generated rationales for the final answers (orange rectangle).
@@ 4.1 Medical Editing Benchmark Construction
Due to the lack of existing benchmarks for medical knowledge editing, we construct two datasets: MedExQAedit and MedMCQAedit by extending two real-world medical QA datasets (Kim et al., 2024; Pal et al., 2022). Each QA dataset is originally provided with a set of questions $q$, answers $a$, and relevant references $c$ in the format of human-written explanations or textbook references.
To ensure reliable evaluations of a target LLM, we filter out medical questions that the LLM can answer correctly before editing. In this way, our assessments could reflect genuine improvements from knowledge editing methods. We prompt the LLM in a zero-shot manner to predict answers for each candidate question and retain only those questions it fails to answer correctly as $\mathcal{Q}_{\mathrm{ori}}$. Thus, the accuracy on $\mathcal{Q}_{\mathrm{ori}}$ indicates the efficacy of editing methods in injecting new knowledge.
To assess the generalization of edited knowledge, we create a novel scenario-based question set $\mathcal{Q}_{\mathrm{gen}}$ that evaluates whether the edited LLM can apply the injected knowledge to address clinical questions in unseen scenarios. Unlike existing benchmarks that rely on simple question paraphrasing and expansions (Ma et al., 2024b; Huang et al., 2024), we generate $\mathcal{Q}_{\mathrm{gen}}$ by extending the core facts of each original question into novel clinical scenarios, and curate $\mathcal{Q}_{\mathrm{ret}}$ by formulating queries that shift focus away from the original topic to assess retention of unaffected knowledge. See Appendix A for data construction details. The examples of $\mathcal{Q}_{\mathrm{ori}}$, $\mathcal{Q}_{\mathrm{gen}}$, and $\mathcal{Q}_{\mathrm{ret}}$ are shown in Figure 9.
@@
As shown in the top gray rectangle of Figure 2, our benchmarks are built following two key data principles:
  • Quality. We filter each QA pair by verifying whether the provided knowledge logically entails the ground-truth answer to ensure the question is knowledge-relevant. We simply prompt the LLM with $(c,q)$ and retain the sample only if the output answer equals $a$.
  • Difficulty. To strictly evaluate the efficacy and generalization of knowledge editing, we ensure the pre-edit performance of the LLM is 0% on $\mathcal{Q}_{\mathrm{ori}}$ and $\mathcal{Q}_{\mathrm{gen}}$, while 100% on $\mathcal{Q}_{\mathrm{ret}}$ to ensure the retention of original knowledge.
4.2 Knowledge-driven Editing Paradigms
@@
Ground-Truth Answer Editing (GTA-Edit) is the prevailing paradigm that typically uses the final answer $a$ as the target knowledge $k$ for editing (Meng et al., 2022, ). As shown in Figure 2, GTA-Edit directly injects the answer (e.g., “Cervical vertebrae”) into LLMs. Although this approach is straightforward, it may lead to superficial memorization of the answer without a deep understanding of the underlying medical rationale.
Reference Editing (RE-Edit) takes supporting reference $c$ extracted from textbooks or academic literature as the target knowledge $k$ for editing. While the reference text is not directly used as the answer, it provides a more comprehensive context for the model to understand the underlying medical knowledge.
Self-generated Rationale Editing (SGR-Edit) is proposed to further enhance the knowledge internalization by leveraging the model’s own reasoning. As shown in the green part of Figure 2, we first prompt the LLM to generate a chain-of-thought rationale for the question-answer pair $(q,a)$ using the reference text $c$ as context. This generated rationale clearly outlines the reasoning process that leads to the answer, serving as the target knowledge $k$ for editing. Thus, SGR-Edit not only incorporates the answer but also provides a detailed explanation of how the model arrives at it, enabling LLMs to deliver interpretable and explainable reasoning that supports high-stakes medical decisions. The detailed prompts for generating rationales are provided in Appendix C.3. Examples of the GTA, RE, and SGR are presented in Table 3.

### v3

@@ Abstract
Knowledge editing (KE) has recently emerged as a promising technique to update specific facts in large language models (LLMs) without full retraining. While existing KE methods show promising results on general-domain benchmarks, their effectiveness in the medical domain remains largely unexplored. Medical knowledge editing poses unique challenges, requiring models not only to memorize new facts but also to internalize and generalize them for reliable and interpretable clinical decision-making. In this work, we propose MedEditBench, a rigorous evaluation framework for assessing medical knowledge editing. Our preliminary results reveal that current KE paradigm, which directly edits simple answers to the LLMs, often leads to superficial updates with poor generalization. To address this, we introduce Self-Generated Rationale Editing (SGR-Edit), which leverages model-generated rationales as editing targets, enabling deeper knowledge integration. Extensive experiments across diverse LLMs and KE methods demonstrate that SGR-Edit consistently improves editing efficacy and generalization. Furthermore, we examine the impact of sequential edits on in-domain medical knowledge, external-domain knowledge, as well as general model capabilities, offering practical insights for deploying KE in real-world medical applications. Codes and datasets are available at https://github.com/Aries-chen/MedEditBench.
@@
Shigeng Chen1thanks:   Equal contribution.,  Linhao Luo211footnotemark: 1,  Zhangchi Qiu1 Yanan Cao3Carl Yang4Shirui Pan1thanks:   Corresponding author. 1School of Information and Communication Technology, Griffith University 2Department of Data Science and AI, Monash University 3Institute of Information Engineering, Chinese Academy of Sciences 4Department of Computer Science, Emory University shigeng.chen@griffithuni.edu.au, s.pan@griffith.edu.au
@@ 1 Introduction
Large language models (LLMs) encapsulate extensive knowledge during training on large-scale corpora (Petroni et al., 2019; Allen-Zhu and Li, 2023; Dai et al., 2021). Nonetheless, their knowledge remains static after training, resulting in factual inconsistencies and hallucinations in tasks that require up-to-date information or domain-specific expertise beyond their pre-trained knowledge (Hu et al., 2023; Huang et al., 2025). Knowledge editing (KE) has emerged as an approach that efficiently updates new knowledge in LLMs, allowing them to constantly adapt to the ever-evolving world without requiring full retraining (Meng et al., 2022, 2023; Fang et al., 2025).
@@
Medical knowledge editing is particularly challenging due to its intricacy and specificity, which requires LLMs to not only memorize updated knowledge (Singhal et al., 2022) but also comprehend underlying concepts and generalize to unseen scenarios (Li et al., 2023), as shown in Figure˜1. Crucially, the necessity for transparent and trustworthy medical decisions requires LLMs to also explain their reasoning to improve trust and interoperability (Holzinger et al., 2017; Tonekaboni et al., 2019). Therefore, these challenges necessitate a more rigorous evaluation framework that accurately reflects the genuine performance of KE methods in the medical domain, ensuring their reliability and applicability in real-world applications (Zhou et al., 2024; Zhang et al., 2025).
In this work, we create a stringent medical knowledge editing benchmark based on two professional medical entrance examinations (Kim et al., 2024; Pal et al., 2022). Specifically, we carefully curate a subset of QA pairs that current LLMs fail to answer correctly to serve as the target set for knowledge editing. To evaluate the generalizability and robustness of KE methods, we extend the benchmark with scenario-based questions that require LLMs to apply the injected knowledge to address new clinical scenarios, as well as preservation questions that require LLMs to retain their previously factual knowledge.
Our findings show that existing KE methods fall short in complex medical settings: edited LLMs typically achieve less than 50% efficacy, struggle to generalize knowledge to new scenarios, and often corrupt previously stored facts. We attribute this pitfall to the existing editing paradigm that edits the LLMs through a simple answer, resulting in surface-level memorization rather than deep understanding, thereby impairing generalization and robustness.
Inspired by research on belief revision in logic and philosophy, where an intelligent agent updates its beliefs by incorporating new information while maintaining logical consistency and minimizing unnecessary changes (Alchourrón et al., 1985; Falappa et al., 2002), we argue that knowledge editing should not merely overwrite outdated facts, but instead be guided by rational principles of consistency and justification. This perspective suggests that, much like human belief revision, effective editing requires encoding not only the updated fact but also the explanatory reasons that support it. Building on this insight, we propose an editing paradigm called Self-Generated Rationale Editing (SGR-Edit), where LLMs are prompted to generate explanatory rationales (Wei et al., 2022) grounded in reference texts. These self-generated rationales serve as the target knowledge that can be seamlessly integrated into arbitrary KE methods. Experimental results demonstrate that SGR-Edit improves knowledge editing performance by up to 11.3% across various LLMs and diverse KE methods, enabling edited LLMs to better internalize new medical knowledge and generalize to unfamiliar scenarios. Our primary contributions are fourfold:
  1. 1.
    The MedEditBench Framework: We introduce MedEditBench, a rigorous evaluation framework for medical knowledge editing. We employ a multi-faceted metric suite to comprehensively assess state-of-the-art KE methods in complex medical scenarios.
  2. 2.
    Pinpointing Paradigmatic Limitations: We conduct extensive evaluations revealing that the conventional KE paradigm, which relies on editing simple final answers, primarily results in superficial memorization rather than genuine knowledge internalization. This severely limits the generalization capability required for real-world practice.
  3. 3.
    Self-Generated Rationale Editing (SGR-Edit): We propose a novel editing paradigm, SGR-Edit, utilizing the LLM’s self-generated, evidence-based reasoning chain as the editing target. It significantly boosts editing performance across diverse LLMs and KE methods.
  4. 4.
    In-depth Analysis and Guidance: We provide the first systematic analysis of sequential editing effects on internal/external knowledge and general capabilities. Our findings offer critical insights for deploying KE in safety-critical domains, guiding future research toward reliable and interpretable LLMs.
@@ 2.1 Knowledge Editing Methods
Current KE methods can be grouped into three main categories. The first, Fine-Tuning–Based Editing, injects new knowledge by updating a large number of model parameters, often through parameter-efficient training techniques (Zhu et al., 2020; Zhang et al., 2024; Hu et al., 2022). The second category, Parameter-Modifying Editing, focuses on altering specific parameters to minimize interference with unrelated knowledge. This group is further divided into: 1) Meta-Learning strategies, which train a hypernetwork to predict gradient updates for knowledge insertion (De Cao et al., 2021; Mitchell et al., 2022a; Tan et al., 2024); and 2) the Locate-then-Edit approach, which directly identifies and rewrites factual weights in specific layers (Meng et al., 2022, 2023), and has recently evolved variants optimized for sequential editing (Fang et al., 2025; Jiang et al., 2025). Finally, Parameter-Preserving Editing maintains the base model’s integrity by either augmenting it with external components (Hartvigsen et al., 2023; Mitchell et al., 2022b; Wang et al., 2024) or using retrieval during inference (Zheng et al., 2023a; Song et al., 2024; Shi et al., 2024; Chen et al., 2024) to store new facts.
Scope of This Work.
While retrieval-based approaches (Song et al., 2024; Shi et al., 2024; Chen et al., 2024) have shown promise by augmenting models with external context, they do not alter the model’s intrinsic knowledge. In this work, we distinguish ourselves by strictly focusing on model editing (parameter updates), aiming to evaluate how well LLMs can intrinsically internalize and generalize medical knowledge without relying on external retrieval modules.
@@ 2.2 Knowledge Editing Benchmarks
Current evaluations of KE methods primarily focused on general-domain benchmarks, such as WikiData (Meng et al., 2022) and counterfactual datasets (Levy et al., 2017), which aimed to update LLMs with new knowledge that contradicted common facts. However, recent studies have questioned the real-world applicability of these methods and proposed new, more rigorous benchmarks for fair evaluations. Specifically, Cohen et al. (2024) measured “ripple effects” on related facts and revealed that most KE methods often fail to propagate consistent changes beyond the target triple. Huang et al. (2024) reported that prior benchmarks do not strictly confirm LLMs having hallucinated answers before editing, which masks their true editing performance. Ma et al. (2024b) investigated editing consistency under prompt rephrasing and realistic contexts, finding that current KE methods exhibit lower generalization, and popular facts are the hardest to edit. Furthermore, Lin et al. (2024) examined sequential editing and showed that KE methods experience performance drops after sequential edits, while Yang et al. (2025b) critiqued common evaluation practices, revealing that KE methods catastrophically fail on realistic QA tasks.
In contrast to previous evaluations focused on simplified general-domain settings (Ma et al., 2024b; Lin et al., 2024), we present the first medical knowledge editing evaluation framework MedEditBench, which not only rigorously measures the effective adoption of injected medical knowledge in new clinical scenarios without degrading previous knowledge but also systematically assesses the effects of sequential edits across internal knowledge, external knowledge, and model general abilities. Furthermore, distinct from recent work focusing on long-tail knowledge Yi et al. (2025), our MedEditBench framework prioritizes the rigorous evaluation of reasoning stability and sequential editing robustness.
@@ 3 Task Formulation
The goal of knowledge editing is to modify specific knowledge $k$ in LLMs without retraining the entire model (Zhang et al., 2024). This modification aims to improve performance on tasks related to that knowledge, represented by a set of queries and answers: $\mathcal{Q}^{k}=\{(q_{i},a_{i})\}$. Let $\theta$ denote the original model. Given a query $q$ and the target knowledge $k$, the knowledge editing method $F$ can be expressed as:
@@ 4 Evaluation Framework
Figure 2: Overview of proposed medical knowledge editing evaluation framework (MedEditBench). Datasets are first constructed for medical knowledge editing (yellow rectangle). Three editing paradigms (GTA-Edit, RE-Edit, and SGR-Edit) are applied to evaluate the effectiveness of different target knowledge for medical knowledge editing (purple rectangle). In particular, the proposed SGR-Edit generates a rationale by prompting the LLM itself, given a QA and a reference (green rectangle). Then, the knowledge is used to edit the base LLM using editing methods like ROME (blue rectangle). In the final evaluation stage (orange rectangle), the edited LLM is evaluated using efficacy, generalization, and retention, as well as interpretability of the generated rationales. To examine how sequential medical edits affect the edited LLM’s broader knowledge and general abilities, we complement MedEditBench with external evaluations, including multi-domain knowledge (MMLU), and commonsense reasoning (HellaSwag), factual recall (TriviaQA), truthfulness (TruthfulQA), and mathematical reasoning (GSM8K).
@@ 4.1 Medical Editing Benchmark Construction
Due to the lack of existing benchmarks for medical knowledge editing, we construct two datasets: MedExQAedit and MedMCQAedit by extending two real-world medical QA datasets (Kim et al., 2024; Pal et al., 2022). Each QA dataset is originally provided with a set of questions $q$, answers $a$, and relevant references $c$ in the format of human‐written explanations or textbook references. To ensure reliable evaluations of a target LLM, we filter out medical questions that the LLM can answer correctly before editing. In this way, our assessments could reflect genuine improvements from knowledge editing methods in the rest set $\mathcal{Q}_{\mathrm{ori}}$. Thus, the accuracy on $\mathcal{Q}_{\mathrm{ori}}$ indicates the efficacy of editing methods in injecting new knowledge. To evaluate generalization and retention, we construct two auxiliary question sets: $\mathcal{Q}_{\mathrm{gen}}$, which extends original facts into new clinical scenarios, and $\mathcal{Q}_{\mathrm{ret}}$, which introduces off-topic questions to assess the preservation of unaffected knowledge. Examples of $\mathcal{Q}_{\mathrm{ori}}$, $\mathcal{Q}_{\mathrm{gen}}$, and $\mathcal{Q}_{\mathrm{ret}}$ are shown in Figure˜9.
@@
As shown in the top gray rectangle of Figure˜2, our benchmarks are built following two key data principles: Quality. We filter each QA pair using Reference Check to verify whether the provided context $c$ logically entails the ground‐truth answer $a$ for the question $q$. We prompt the LLM with $(c,q)$ and retain the sample only if the predicted answer $\hat{a}$ equals $a$. Difficulty. To strictly evaluate the efficacy and generalization of knowledge editing, we ensure the pre-edit performance of the LLM is 0% on $\mathcal{Q}_{\mathrm{ori}}$ and $\mathcal{Q}_{\mathrm{gen}}$ using Zero-Shot Filter (i.e., LLM’s explicit incorrectness serves as direct signal for selecting examples to be targeted for editing and generalization evaluation), while 100% on $\mathcal{Q}_{\mathrm{ret}}$ to ensure the retention of original knowledge. Prompts and details for data construction and verification are provided in Appendix A.2.
4.2 Knowledge‐driven Editing Paradigms
@@
Ground‐Truth Answer Editing (GTA‐Edit) is the prevailing paradigm that typically uses the final answer $a$ as the target knowledge $k$ for editing (Meng et al., 2022, 2023). As shown in Figure˜2, GTA‐Edit directly injects the answer (e.g., “Cervical vertebrae”) into LLMs. Although this approach is straightforward, it may lead to superficial memorization of the answer without a deep understanding of the underlying medical rationale.
Reference Editing (RE‐Edit) takes supporting reference $c$ extracted from textbooks or academic literature as the target knowledge $k$ for editing. It provides a more comprehensive context to understand the underlying new knowledge.
Self‐generated Rationale Editing (SGR‐Edit) is motivated by the rational postulates of belief revision (Alchourrón et al., 1985), which emphasizes that revision must be consistent and closed under logical consequence. Moreover, prior work has argued that intelligent agents typically require an explanation to legitimize newly provided information, and the main role of an explanation is to rationalize facts (Falappa et al., 2002). Guided by these insights, we first prompt the LLM to generate a chain-of-thought rationale for the question-answer pair $(q,a)$ using the reference text $c$ as context. This generated rationale clearly outlines the reasoning process that leads to the answer, serving as the target knowledge $k$ for editing, as shown in the green part of Figure˜2. Thus, SGR-Edit not only incorporates the answer but also provides a detailed explanation of how the model arrives at it, enabling LLMs to deliver interpretable and explainable reasoning that supports the new fact. The detailed prompts for generating rationales are provided in Appendix C.3. Examples of the GTA, RE, and SGR are presented in Table˜3.